I need help finding the zeros of the function algebraically.
f(x)=2x2-7x-30

Answer by stanbon(48558)

f(x)=2x2-7x-30
------------
2x^2-7x-30 = 0
2x^2-12x+5x-30 = 0
2x(x-6)+5(x-6)=0
(x-6)(2x+5) = 0
x = 6 or x = -5/2
=============
Cheers,
Stan H.